# Less for More: Enhanced Feedback-aligned Mixed LLMs for Molecule Caption Generation and Fine-Grained NLI Evaluation

### v2

@@ 4.2.1 Aligning Molecule-Language Modals with Minimal Training
We first present results for molecule language models with minimal alignment fine-tuning, initialising pretrained weights from molecule generation rather than deploying model merging (see Appx. G for details). Tables 2 and 3 summarise experimental results. Generally, benchmarking models trained on extensive data with SFT exhibit memorisation effects, with performance dropping by 50% to 100% compared to reported results, when evaluated on out-of-distribution data.
Our experiments show that not all alignment optimisations are effective in the minimal training setting. Both DPO and KTO show zero performance in caption generation when models are initialised with crossmodal weights unrelated to the task (see Table 2). However, performance improves significantly when the crossmodals are known (see Table 3). In molecule generation, DPO achieves up to 42% better performance than Meditron, trained on the full dataset, while KTO still performs poorly, likely due to overfitting (see Appx. F).
By contrast, CPO effectively handles both the crossmodal agnostic and minimal training settings, outperforming Meditron by up to 20% in caption generation and 42% in molecule generation. This is likely due to its inherent ability to balance structured learning and generalisation. It aligns with preferred data through behaviour cloning and SFT, which encourage the model to mimic expert behaviour, while reducing bias and suboptimal outcomes via a uniform reference model that assigns equal likelihood to all possible outputs.
@@
4.2.2 Synergies Between Model Merging and Alignment Fine-tuning
@@
Overall, our experiments show that model merging can effectively address key limitations in alignment fine-tuning. By fusing pretrained models, one can enhance performance with minimal training, reducing reliance on human-labelled data, lowering training costs, minimising human bias, and improving generalisation. Examples of caption and molecule generation are provided in Appx. H.
@@ 4.2.3 Atomic-level Cross-NLI Evaluation
Atomic-level NLI revealed intriguing insights regarding performance interpretation. Fig. 4 shows assessment score distributions from our proposed evaluation method, comparing our top models against Meditron trained on the entire dataset. All models exhibit low hallucination, likely due to the narrow, well-defined topics that enable factually correct captions without unrelated information. However, our models excel in coverage, generating more comprehensive captions, with performance increasing to 69% compared to Meditron’s 51% (Fig. 4 (B)). Examples of insights captured by our proposed evaluation are in Appx. I.
@@
We compare our atomic-level NLI method against established approaches, including full NLI, which treats entire premises and hypotheses as single units, and sentence-level NLI Laban et al. (2022)), which evaluates smaller parts. Using outputs from our top-performing model, CPO+SLERP, as preferred captions and those from Meditron as dispreferred ones, we measure the relative entropy in coverage score distributions. The results indicate that both full and sentence-level NLI methods misclassify preferred captions as non-entailment and dispreferred ones as entailment (Fig. 5 (B)-(D)). By contrast, atomic-level cross-NLI accurately favours preferred captions, assigning higher scores to certain cases (Fig. 5 (A)). Additionally, Kullback–Leibler divergence shows that atomic-level NLI offers better discrimination, achieving a divergence score of 0.54 compared to 0.12–0.17 for other methods, demonstrating its effectiveness in distinguishing the quality of generated captions.
@@ 5 Conclusion
In this work, we address limitations of scientific language models that rely on extensive training. Focusing on molecule caption generation, we propose synergies between model merging and alignment fine-tuning with minimal training to enhance chemical language models. Our experiments show that while alignment fine-tuning performs poorly, incorporating model merging significantly outperforms extensively trained models on out-of-distribution data, offering a cost-effective approach that relies less on human-labelled data. Furthermore, we propose an atomic-level cross-NLI evaluation to overcome limitations of widely used NLI evaluation methods, which lack appropriate granularity. Our method provides valuable insight into performance interpretability and effectively handles multiple content units, where existing NLI methods consistently misalign with assessment criteria.
@@
In this work, we employ weight-based and subspace-based merging methods to create universal models for the MoCG task, facilitating alignment fine-tuning in a training setting with minimal data. However, both are static merging methods. This means that the merged model remain the same for all samples or tasks. Given that there are differences between input samples/tasks, the models’ ability may vary when processing different samples/tasks. In the future, we aim to investigate dynamically merging models (or subsets of layers) based on the samples/tasks during the inference phase Kang et al. (2024).
We also propose an atomic-level NLI evaluation method that successfully handles multiple content units, offering valuable insights into performance interpretability for caption generation, where widely adopted NLI methods consistently misalign with assessment criteria. However, decomposing text into atomic units can be challenging for other tasks involving complex or lengthy text. While this method captures nuanced content, there is a risk of over-fragmentation, which may lead to a loss of context or coherence in evaluation. Additionally, the effectiveness of this approach relies heavily on the LLM for decomposition and the NLI model for entailment and contradiction assessment. If either model struggles with domain-specific content (e.g., highly technical language), the evaluation could yield inaccurate or biased results. Furthermore, if generated texts introduce valid but creative or non-standard content, this approach may penalise them by classifying such deviations as contradictions or hallucinations, even when they provide accurate information. Future work will need to address these limitations across various domains.
@@
Appendix A Foundations in Alignment with RLHF
@@ Appendix B Data
Appendix C LLM Instruction for Decomposing Text into Atomic Context Units
@@
Appendix D Instructions for Molecule Language Translation
@@
Appendix E Evaluation Metrics
@@
Appendix F Training Efficiency
@@
Appendix G Implementation Details
All implementations used Meditron Chen et al. (2023) as the backbone model, known for its performance on L+M-24. For alignment fine-tuning experiments, we initialised Meditron crossmodals, trained for molecule generation  [^4] Crossmodal initialisation was based on the most challenging task reported in Edwards et al. (2024).. For the model merging experiments, we combined Meditron weights trained on MoCG tasks in a 19:1 ratio. This ratio aimed to preserve the balance of information between the linguistic and molecule modalities. All models were fine-tuned using QLoRA Dettmers et al. (2024).
@@ Appendix G Implementation Details
# Training Configurations args = TrainingArguments( output_dir=save_path, overwrite_output_dir=True, load_best_model_at_end=True, num_train_epochs=3, per_device_train_batch_size=1 per_device_eval_batch_size=1 gradient_accumulation_steps=64 gradient_checkpointing=False optim="adamw_torch_fused", learning_rate=5e-5, max_grad_norm=0.3, warmup_ratio=0.1, lr_scheduler_type="cosine", ) # Quantisation Configurations ( load_in_4bit=True, bnb_4bit_use_double_quant=True, bnb_4bit_quant_type=nf64, bnb_4bit_compute_dtype=torch.bfloat16 ) # LoRA Configurations ( lora_alpha=16, r = 64, lora_dropout=0.1, task_type="CAUSAL_LM", bias=False, target_modules= "all-linear" )
@@
Appendix H Examples of generated molecules and captions.
Fig. 9 and 10 illustrate examples of molecules and captions generated by our top-performing models compared to Meditron, respectively.
@@
Appendix I Examples of Atomic-level Cross-NLI evaluation

### v3

@@ 4.2.1 Aligning Molecule-Language Modals with Minimal Training
We first present results for molecule language models with minimal alignment fine-tuning, initialising pretrained weights from molecule generation rather than deploying model merging (see Appx. J for details). Tables 2 and 3 summarise experimental results. Generally, benchmarking models trained on extensive data with SFT exhibit memorisation effects, with performance dropping by 50% to 100% compared to reported results, when evaluated on out-of-distribution data.
Our experiments show that not all alignment optimisations are effective in the minimal training setting. Both DPO and KTO show zero performance in caption generation when models are initialised with crossmodal weights unrelated to the task (see Table 2). However, performance improves significantly when the crossmodals are known (see Table 3). In molecule generation, DPO achieves up to 42% better performance than Meditron, trained on the full dataset, while KTO still performs poorly, likely due to overfitting (see Appx. I).
By contrast, CPO effectively handles both the crossmodal agnostic and minimal training settings, outperforming Meditron by up to 20% in caption generation and 42% in molecule generation. This is likely due to its inherent ability to balance structured learning and generalisation. It aligns with preferred data through behaviour cloning and SFT, which encourage the model to mimic expert behaviour while reducing bias and suboptimal outcomes via a uniform reference model that assigns equal likelihood to all possible outputs.
@@
4.2.2 Alignment with Model Merging
@@
For our best-performing model, CPO+SLERP, we conducted ablation studies to assess the impact of weight interpolation coefficients when merging pretrained models on MoCG tasks. Specifically, we explored blending weights across all layers (0–32) to preserve Mol2Cap performance while improving Cap2Mol performance (see Appx. A for details), aiming to create a universal model with enhanced overall capability. Fig. 4 shows the performance trends across different mixing ratios of per-task model weights. Empirically, we found that a 1:18 ratio (Mol2Cap:Cap2Mol) yields the best balance, favoring Mol2Cap performance at lower ratios and Cap2Mol performance at higher ones. Further comparison with a baseline method, namely model soup Wortsman et al. (2022), is provided in Appx. A.
Figure 4: Ablation of best performance model, CPO+SLERP, for Mol2Cap and Cap2Mol tasks, evaluating the effect of per-task model weight mixing ratios.
Overall, our experiments show that model merging mitigates key limitations in alignment fine-tuning. By fusing pretrained models, it boosts performance with minimal training, reduces reliance on human-labeled data, lowers costs, minimizes bias, and improves generalization. Caption and molecule generation examples are in Appx. K.
@@ 4.2.3 Atomic-level Cross-NLI Evaluation
Atomic-level NLI revealed intriguing insights regarding performance interpretation. Fig. 5 shows assessment score distributions from our proposed evaluation method, comparing our top models against Meditron trained on the entire dataset. All models exhibit low hallucination, likely due to the narrow, well-defined topics that enable factually correct captions without unrelated information. However, our models excel in coverage, generating more comprehensive captions, with performance increasing to 69% compared to Meditron’s 51% (Fig. 5 (B)). Examples of insights captured by our proposed evaluation are in Appx. L.
@@
We also evaluated the robustness of our proposed NLI evaluation method against leading approaches by measuring the relative entropy of textual entailment between human-curated texts (i.e., gold labels) and outputs generated by our top-performing model, CPO+SLERP (preferred), versus those from a low-performing model, Meditron (dispreferred). Ideally, all NLI methods should favour preferred outputs over dispreferred ones. However, we observed that both the full and sentence-level NLI methods misclassify preferred captions as non-entailment and dispreferred captions as entailment (see Fig. 6 (B)-(D)). By contrast, atomic-level cross-NLI accurately favours preferred captions, assigning higher scores to certain cases (Fig. 6 (A)). Additionally, Kullback–Leibler divergence shows that atomic-level NLI offers better discrimination, achieving a divergence score of 0.54 compared to 0.12–0.17 for other methods, demonstrating its effectiveness in distinguishing the quality of generated captions. We leave further ablation analysis in Appx. B.
@@
We conducted ablation studies on our atomic-level NLI evaluation method to assess its semantic robustness, particularly in handling complex, lengthy captions that may lose cohesiveness due to excessive decomposition into atomic units. To evaluate this, we analyzed the word count distribution (see Appx. B), filtered captions with at least 50 words, and recalculated relative entropy against standard NLI methods. Fig. 7 demonstrates the superior performance of our method on longer cases compared to leading alternatives. Our NLI method demonstrated a significant improvement in its ability to differentiate preferred outputs from dispreferred ones accurately, achieving a KL divergence of 2.53 (see Fig. 7), as opposed to a KL divergence of 0.54 across all cases in the test subset (see Fig. 6). In contrast, other leading NLI methods experienced a marked increase in KL divergence, favouring dispreferred outputs, which misaligned with the entailment aspect.
Figure 7: Relative entropy in coverage scores for preferred vs. dispreferred generated captions across atomic-level and leading NLI approaches in long captions.
@@ 5 Conclusion
In this work, we address limitations of scientific language models that rely on extensive training. Focusing on molecule caption generation, we propose synergies between model merging and alignment fine-tuning with minimal post-training to enhance chemical language models. Our experiments show that while alignment fine-tuning performs poorly, incorporating model merging significantly outperforms extensively trained models on out-of-distribution data, offering a cost-effective approach that relies less on human-labelled data. Furthermore, we propose an atomic-level cross-NLI evaluation to overcome limitations of widely used NLI evaluation methods, which lack appropriate granularity. Our method provides valuable insight into performance interpretability and effectively handles multiple content units, where existing NLI methods consistently misalign with assessment criteria.
@@
In this work, we employ weight-based and subspace-based merging methods to create universal models for the MoCG task, facilitating alignment fine-tuning in a training setting with minimal data. However, both are static merging methods. This means the merged model remains the same for all samples or tasks. Given that there are differences between input samples/tasks, the models’ ability may vary when processing different samples/tasks. In the future, we aim to investigate dynamically merging models (or subsets of layers) based on the samples/tasks during the inference phase Kang et al. (2024); Yang et al. (2024).
We also propose an atomic-level NLI evaluation method that successfully handles multiple content units, offering valuable insights into performance interpretability for caption generation, where widely adopted NLI methods consistently misalign with assessment criteria. However, decomposing text into atomic units can be challenging for other tasks involving complex or lengthy text. While this method captures nuanced content, there is a risk of over-fragmentation, which may lead to a loss of context or coherence in evaluation. Additionally, the effectiveness of this approach relies heavily on the LLM for decomposition and the NLI model for entailment and contradiction assessment. The evaluation could yield inaccurate or biased results if either model struggles with domain-specific content (e.g., highly technical language). Furthermore, if generated texts introduce valid but creative or non-standard content, this approach may penalise them by classifying such deviations as contradictions or hallucinations, even when they provide accurate information. Future work will need to address these limitations across various domains.
@@
Acknowledgements
This work was supported by a UKRI/EPSRC Turing AI Fellowship to Maria Liakata (grant ref EP/V030302/1) and the Alan Turing Institute (grant ref EP/N510129/1). This work was supported by the Engineering and Physical Sciences Research Council [grant number EP/Y009800/1], through funding from Responsible Ai UK (KP0016) as a Keystone project lead by Maria Liakata.
Appendix A Complementary Experiments in Model Merging
We compared SLERP and TIER model merging techniques against a weighted linear combination of parameters, referred to as model soup Wortsman et al. (2022), when applying CPO in the MoCG task. Our results indicated that model soup caused a significant drop in performance for both Mol2Cap and Cap2Mol tasks (see Fig. 8). We hypothesise that this is because model soup assumes that performance improvement or preservation is linearly related to weight blending, which may not hold for complex models. This observation justifies our decision to explore task-specific arithmetic and geometric merging approaches, as they inherently manage conflicts and better preserve the strengths of each model in specialised tasks.
Figure 8: Comparison of SLERP and TIES with Model Soup for (A) Mol2Cap and (B) Cap2Mol generation.
Appendix B Complementary Experiments in Our Atomic-Level NLI Evaluation Method
First, we analysed the distribution of word counts in captions from the test subset. We observed that the captions are typically short, with an average of 31 words (STD = 50) as shown in Fig.9. Additionally, the captions generally exhibit little dependency across sentences, as they consist of simple natural language describing chemical properties (for a more detailed view, see Table 6).
Figure 9: Distribution of word counts in captions from the test subset.
[…]
Figure 10: Relative entropy in coverage scores for preferred vs. dispreferred generated captions across atomic-level and leading NLI approaches in extreme captions.
Appendix C Foundations in Alignment with RLHF
@@ Appendix D Data
Appendix E LLM Instruction for Decomposing Text into Atomic Context Units
@@
Appendix F Instructions for Molecule Language Translation
@@
Appendix G Baselines
  • TxtChem-T5 Christofidellis et al. (2023) is a T5XL multitask model trained on linguistic and molecule modalities across multiple datasets, including CheBI-20, akin to L+M-24.
  • Chem-LLM Zhang et al. (2024), an InternLM2-Base-7B model, is trained on large chemical knowledge databases using DPO, achieving GPT-4-level results.
  • Meditron Chen et al. (2023), a 7B model, is fine-tuned on the entire L+M-24 dataset.
Appendix H Evaluation Metrics
@@
Appendix I Training Efficiency
@@
Appendix J Implementation Details
All implementations used Meditron Chen et al. (2023) as the backbone model, known for its performance on L+M-24. For alignment fine-tuning experiments, we initialised Meditron crossmodals, trained for molecule generation  [^4] Crossmodal initialisation was based on the most challenging task reported in Edwards et al. (2024).. For the model merging experiments, we combined Meditron weights trained on MoCG tasks in a 1:18 ratio. This ratio aimed to preserve the balance of information between the linguistic and molecule modalities. All models were fine-tuned using QLoRA Dettmers et al. (2024).
@@
( load_in_4bit=True, bnb_4bit_use_double_quant=True, bnb_4bit_quant_type=nf64, bnb_4bit_compute_dtype=torch.bfloat16 )
Figure 13: Quantisation Configurations
args = TrainingArguments( output_dir=save_path, overwrite_output_dir=True, load_best_model_at_end=True, num_train_epochs=3, per_device_train_batch_size=1 per_device_eval_batch_size=1 gradient_accumulation_steps=64 gradient_checkpointing=False optim="adamw_torch_fused", learning_rate=5e-5, max_grad_norm=0.3, warmup_ratio=0.1, lr_scheduler_type="cosine", )
@@
( lora_alpha=16, r = 64, lora_dropout=0.1, task_type="CAUSAL_LM", bias=False, target_modules= "all-linear" )
Figure 15: LoRA Configurations
Appendix K Examples of generated molecules and captions.
Fig. 16 and 17 illustrate examples of molecules and captions generated by our top-performing models compared to Meditron, respectively.
@@
Appendix L Examples of Atomic-level Cross-NLI evaluation